What is the percent ionization in a 0.00350 M solution of base that has a pH of 8.51?

Report your answer in the correct number of significant figures. Report the percentage, not the percentage as a decimal value. For example, if you calculate 50.0 %, report 50.0, not 0.500.
Answer #1


POH=14.00- pH =14.00 -8.51 = 5.49 [OH-] =10-549 =3.24x10-6 [OH-] % of ionization == Initial concentration _ 3.24 x 106 0.0035